solve the equation
5. x2 – 3x + 2 = 0

Answer by ewatrrr(10682)


Hi,
solving the equation
x^2 – 3x + 2 = 0
factoring
(x-1)(x-2)= 0 Note:SUM of the inner product(-1*x) and the outer product(-2x) = -3x
(x-1)=0
x = 1
(x-2)= 0
x = 2